OpenStudy (anonymous):

how about this one: \[(3-6i)^{2}\]

OpenStudy (anonymous):

-27 -36i

OpenStudy (anonymous):

45 - 36i

OpenStudy (anonymous):

Second one is right :D

OpenStudy (anonymous):

you have a choice but the easiest thing is just multiply out \[(3-6i)(3-6i)=9-36i+36i^2=9-36i-36=-27-36i\]

OpenStudy (anonymous):

Lol, I got it right the first time. Thanks satellite :D

OpenStudy (anonymous):

think you missed this daniel or else i am very tired. (also true)

OpenStudy (anonymous):

No, I missed it. I second guessed myself and came up stupidly :D

OpenStudy (anonymous):

oh i didn't see first one was you as well. sorry.

OpenStudy (anonymous):

I was wrong. You were right. Simple as that :D Good catch satellite :D